OpenStudy (anonymous):

Verify that the function satisfies the three hypotheses of Rolle's Theorem on the given interval. Then find all numbers c that satisfy the conclusion of Rolle's Theorem. (Enter your answers as a comma-separated list.) f(x) = x3 − x2 − 2x + 5, [0, 2]

OpenStudy (anonymous):

Derivative is 3x^2-2x-2 F(2) = 5 F(0) = 5 F(2)=F(0) = 0 3x^2-2x-2=0